2. A tank, initially empty, is in the form of a right circular cylinder with hemispherical ends. The overall

length of the cylinder tank is 3 meters and the diameter of the hemisphere is 1 meter. If a pump

discharges a fluid, whose density is 1.2 kgm/Liter, in this tank at a rate of 300 liters per minute

determine a.) Weight of liquid inside the tank after 5 minutes, KN, and b) total time to fill the tank,

min.


1
Expert's answer
2022-02-21T14:52:56-0500

Volume of hemisphere

"V=\frac{1}{2}\frac{4}{3}\pi\frac{d^3}{8}=\frac{\pi d^3}{12}"

Volume of the right circular cylinder

"V_2=\pi\frac{d^2}{4}(l-2\times\frac{d}{2})=\frac{\pi d^2}{8}(l-d)"

"M=\rho(V_1+V_2)"

Put value

"M=550\pi kg m"

"M=1728kg"

Find the mass of liquid inside the tank if it is half full.

"W=mg"


"W=550\times3.14\times9.8=16950N"

"V=\frac{\pi d^2}{6}"

Volume

"V=2V_1+V_2"

Put value

"V=\frac{13\pi}{24}m^3"

Now time

"t=\frac{\frac{13\pi}{24}}{0.3}=5.6min"